# Grasping a Handful: Sequential Multi-Object Dexterous Grasp Generation

### v2

@@ I Introduction
Generation of dexterous grasps has been studied for a long time, both from a technical perspective on generating grasps on robots [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and understanding human grasping [12, 13, 14, 15]. Most of these methods rely on bringing the robotic hand close to the object and then simultaneously enveloping it with all fingers. While this strategy often results in efficient and successful grasp generation, it simplifies dexterous grasping to resemble parallel-jaw grasping, thereby underutilizing the many DoF of multi-fingered robotic hands [10]. In contrast, grasping multiple objects with a robotic hand, particularly in a sequential manner that mirrors human-like dexterity, as shown in Fig. 1, is still an unsolved problem.
@@
  • SeqGrasp, a novel hand-agnostic algorithm for sequential multi-object grasp generation.
@@ II-B Multi-Object Grasping
A few works have addressed dexterous multi-object grasping [16, 25, 28] where [16] targets simultaneous multi-object grasping while [25, 28] targets sequential multi-object grasping. Li et al. [16] proposed MultiGrasp a two-stage simultaneous multi-object dexterous grasping framework where a generative grasp sampler proposed poses to simultaneously pick many objects, followed by a learned policy for executing the pick. The main limitation of [16] is that objects must be spatially close and of similar size and shape. In comparison, our method can handle scattered objects of different shapes and sizes by sequentially picking one at a time. The other works that do sequential multi-object grasping [25, 28] restrict the grasping to a maximum of two objects [25] or to primitive object shapes such as cylinders or spheres [28]. In comparison, our method can handle up to four objects of complex shapes and sizes. We also use our method to collect the largest sequential multi-object grasping dataset to date.
@@ III Problem Formulation
$\mathbf{g}_{n}=\operatornamewithlimits{argmin}_{\mathbf{g}_{n}}E(\mathbf{g}_{n},\mathbf{O}_{n},\mathcal{G}_{n-1},$ $\;\mathcal{O}_{n-1},\mathcal{OS}_{n}),$
$\;\quad\forall n=1,\dots,N,$ (1)
where $\mathcal{G}_{n-1}=\{\mathbf{g}_{i}\}_{i=1}^{n-1}$, $\mathcal{O}_{n-1}=\{\mathbf{O}_{i}\}_{i=1}^{n-1}$, $\mathcal{G}_{0}=\varnothing$, and $\mathcal{O}_{0}=\varnothing$. $E$ in Eq. 1 is a differentiable function that quantifies how well grasp $\mathbf{g}_{n}$ can pick object $\mathbf{O}_{n}$ with the DoF $\mathcal{OS}_{n}$ given all previously generated grasps $\mathcal{G}_{n-1}$ and objects $\mathcal{O}_{n-1}$.
@@ IV-A Opposition Space Selection Strategy
The primary objective in sequential multi-object grasping is to maximize the hand’s remaining DoF after each grasp. For this purpose, we propose a grasp planning strategy guided by OSes [33, 14, 34, 28]. An OS is a functional subspace within the hand’s kinematic structure formed by pairs of opposing surfaces (such as fingertips, lateral surfaces of fingers, or palm surfaces) along with the joints that control these surfaces [28]. It represents regions where opposing forces can be applied to create stable grasps. The number of OSes is hand-dependent and varies based on the kinematic structure. Fig. 3(a) shows the seven different OSes for the Allegro Hand.
Mathematically, each opposition space can be represented as a pair $\mathcal{OS}_{i}=\{\mathbf{J}_{i},\mathbf{S}_{i}\}$, where $\mathbf{J}_{i}\in\{0,1\}^{K}$ is a binary vector indicating which joints are involved in controlling the opposition space, and $\mathbf{S}_{i}\in\mathbb{R}^{3\times M_{i}}$ represents the 3D points on the hand where opposing forces can be applied. Fig. 3(b) shown an example of two different $\mathbf{S}_{i}$ for the Allegro Hand, where palm and pad oppositions have contact points located on the inner surfaces of fingers and palm and side oppositions have contact points on the fingers’ lateral surfaces.
@@ IV-B Optimization-based Grasp Generation
We initialize $\mathbf{g}_{n}$ at a randomly sampled position on the expanded convex hull of the target object $\mathbf{O}_{n}$ as exemplified in Fig. 4(a),. If $n=1$, then $\bm{\theta}_{1}$ is set to a natural open-hand and collision-free posture, while for $n\geq 2$, $\bm{\theta}_{n}=\bm{\theta}_{n-1}$. A visual example of the optimization process when grasping a second object is shown in Fig. 4(b).
@@ IV-C Energy Function
The force-closure term ($E_{\text{fc}}$) encourages the grasp to be in force-closure equilibrium [36]. Following [10] and assuming zero friction and uniform contact force magnitudes, we define it as
@@ V Dataset Generation
max width= OSes Success Total Success Rate (%) Middle-Ring 100.83 704.97 14.30 Index-Middle 97.43 700.90 13.90 Thumb-Index 97.29 1691.41 5.75 Ring-Palm 147.33 323.31 45.57 Middle-Palm 152.77 404.84 37.74 Index-Palm 194.13 546.60 35.52 Thumb-Palm 85.86 485.80 17.67 Num. Obj. Grasped Consumed Total Consumed Rate (%) One 0.00 444.72 0.00 Two 9.75 251.52 3.88 Three 69.20 139.31 49.68 Four 40.09 40.09 100.00
@@ VII Experiments
Refer to captionSeqGraspSeqDiffuserMultiGraspSeqDatasetGrasp’Em
Figure 5: Qualitative results. For SeqGrasp and SeqDiffuser, we only show consumed grasps, that is, when $\mathcal{OS}=\emptyset$. For SeqGrasp and SeqDiffuser, the grasp sequences are visually indicated by a color gradient, transitioning from lighter to darker shades. In contrast, for MultiGrasp, the color gradient is only used to differentiate the objects.
We experimentally evaluate SeqGrasp and SeqDiffuser in both simulation and the real world. The specific questions we wanted to address with the experiments were:
@@
We compare our method to the optimization-based sampler MultiGrasp from [16], which generates simultaneous grasps on clustered objects. As such, for MultiGrasp, we must first sample clustered object configurations and then generate multi-object grasps directly on the object cluster. In contrast, SeqGrasp and SeqDiffuser do not require objects to be spatially close as they generate grasps sequentially based on previously successful ones. While this comparison is not entirely fair, we still believe comparing these two strategies offers valuable insights.
@@ VII-A Simulation Experiments
    Diversity (Div.) in radian: Grasp diversity is determined by calculating the standard deviation of $\mathbf{g}$ across all successful grasps.
@@
    Efficiency (Eff.) in second: The computational time required to generate a batch of 256 grasps on an NVIDIA A100.
@@
max width= SR $\uparrow$ Pene. $\downarrow$ Eff. $\downarrow$ Div. $\uparrow$ Method SData G’Em SData G’Em Avg. Avg. MulG-1 66.84 65.39 1.14 1.27 600 0.284 SeqD-1 46.95 45.23 5.55 5.59 0.8 0.321 SeqG-1 50.04 40.78 1.73 2.16 900 0.332 MulG-2 22.46 16.48 2.30 2.83 750 0.347 SeqD-2 18.83 23.00 5.84 6.28 0.8 0.359 SeqG-2 21.21 32.03 2.14 1.78 900 0.367 MulG-3 10.78 3.55 3.39 4.04 900 0.340 SeqD-3 11.05 9.22 5.93 6.47 0.8 0.334 SeqG-3 19.49 21.05 2.23 2.23 900 0.349 MulG-4 0.90 0.47 5.17 6.27 1000 0.329 SeqD-4 3.01 1.68 6.18 6.69 0.8 0.293 SeqG-4 2.93 5.04 2.70 2.62 900 0.312
TABLE III: Simulation results. MulG, SeqG, SeqD, G’Em, and SData are short the MultiGrasp, SeqGrasp, SeqDiffuser, Grasp’Em, and SeqDataset, respectively. The $-i$ following the method name denotes the number of objects used for grasp generation. $\uparrow$($\downarrow$), the higher (lower), the better.
The quantitative results are presented in Table III, while Fig. 5 qualitatively illustrates a few grasps. The results demonstrate that SeqGrasp achieves the highest success rate and the lowest penetration depth when grasping two or more objects. MultiGrasp performs well for one- and two-object grasps, as it utilizes all of the hand’s available DoF to grasp the objects. However, because MultiGrasp requires all objects to be initialized nearby, the success rate of the generated grasps is susceptible to the initial object placements. In contrast, SeqGrasp and SeqDiffuser do not suffer from this limitation.
We observe a significant performance drop when transitioning from three-object to four-object grasps across all methods. We hypothesize that this decline occurs because the three previously grasped objects occupy substantial space within the Allegro Hand, pushing the fourth object grasp to the limits of the hand’s kinematic redundancy. Additionally, as the number of grasped objects increases, object-object interactions grow exponentially, making the task considerably more challenging, a finding also reported in [16]. Nevertheless, SeqGrasp demonstrates superior performance in scenarios involving three or more objects.
Another notable observation is that SeqDiffuser generates grasps with high penetration, which aligns with previous work on single object diffusion-based grasp sampling [5]. Still, SeqDiffuser is valuable because it generates grasps 750-1250 times faster than SeqGrasp and MultiGrasp.
@@ VII-B Real-World Experiments
To address the final question, we evaluated the stability of the generated grasps on a real Allegro Hand mounted to a Franka Emika Panda robotic arm. For the evaluation, we 3D printed the same eight SeqDataset test objects from the simulation experiments. The printed objects are shown in Fig. 6. We applied a silicone gel coating to the object surfaces and attached an anti-slip grip to the palm to increase friction.
We followed the procedure outlined in [28] to replicate the grasp on the real hardware. This procedure involves positioning each object as closely as possible with its generated pose and then closing the relevant joints of the hand to retain the objects. For grasps with severe penetration issues, we placed the object in the closest feasible real-world position. We only evaluated three- and four-object grasps. For MultiGrasp and SeqGrasp, we replicated the successful grasps with the lowest energy, while for SeqDiffuser, we randomly chose one of the generated grasps.
We evaluated the grasps with the grasp stability test shown in Fig. 6. In this test, the arm was first moved to a holding position with the palm facing downward. Next, the arm moved left-to-right, and then the last joint rotated $\pm$90°. A trial is considered successful if all objects remain grasped.
Figure 6: Left: Printed object set with silicone gel coating over the surface. Right: The stability test.
@@
max width= Real exp. SeqG-3 SeqG-4 SeqD-3 SeqD-4 MulG-3 MulG-4 Succ. trials 6/10 6/10 2/10 1/10 2/10 0/7 Succ. objects 24/30 33/40 20/30 21/40 11/30 12/28
TABLE IV: Real-world experimental results. Succ. trials mean successful trials and succ. objects mean how many objects were still grasped even if the trial failed.
The experimental results are presented in Table IV. Similar to the simulation results, SeqGrasp outperformed SeqDiffuser and MultiGrasp in both three-object and four-object grasping tasks, achieving an average success rate of 60%. The primary factor contributing to SeqDiffuser’s low success rate (15%) was significant object interpenetration in the generated grasps, which meant that the closest physically feasible grasp replicable on the real hand differed substantially from the intended configuration. MultiGrasp, with an average success rate of (12%), mainly failed because the stability of the generated grasps relied on many object-object contacts, which are sensitive to minor object displacements. This resulted in frequent failures, particularly in four-object grasping trials where MultiGrasp achieved no successful trials and was even unsuccessful in finding solutions in three trials, which is why the last column in Table IV is seven and not ten. If we also count these as failed trials, the average success rate for MultiGrasp decreases to 10%.
Our real-world experiments highlight the robustness of sequential grasping as a strategy for dexterous multi-object manipulation. By generating grasps iteratively and independently of object proximity, SeqGrasp and SeqDiffuser mitigate the challenges posed by object interactions, making it a more reliable and practical approach for real-world applications.
VIII Limitations
Although both SeqGrasp and SeqDiffuser can produce high-quality and diverse grasps, the generated grasps represent only a subset of possible multi-object grasp solutions. Specifically, our solution relies on only one or two fingers for a grasp. In contrast, humans can flexibly adjust their fingers and distribute space within the hand to grasp objects of varying sizes, shapes, and weights. Moreover, our approach does not account for object sequencing, whereas humans may prefer a specific order to facilitate multi-object grasping. Addressing these limitations is an exciting future research direction.

### v3

@@ I Introduction
Dexterous grasping has been extensively studied in robotics [1, 2]. The field has evolved through two primary research directions. First, numerous methods have been developed for generating robotic grasps [3, 4, 5, 6, 7, 8], from analytical approaches [7] to data-driven generative models including Generative Adversarial Networks [5], diffusion models [4, 3], and Variational Autoencoders [8, 6]. Second, a complementary line of research has focused on understanding human grasping behaviors, including intention recognition [9], real-time tracking of hand-object interactions [10], grasp taxonomy classification [11], and learning task-oriented grasping strategies from human demonstrations [12]. The primary strategy in dexterous grasp generation has been to bring the robotic hand close to the object and then simultaneously envelop it with all fingers. While this strategy often results in efficient and successful grasp generation, it simplifies dexterous grasping to resemble parallel-jaw grasping, thereby underutilizing the many DoF of multi-fingered robotic hands [7].
@@
  • SeqGrasp, a novel algorithm for sequential multi-object grasp generation.
@@ II-B Multi-Object Grasping
A few works have addressed dexterous multi-object grasping [13, 22, 25] where [13] targets simultaneous multi-object grasping while [22, 25] targets sequential multi-object grasping. Li et al. [13] proposed MultiGrasp a two-stage simultaneous multi-object dexterous grasping framework where a generative grasp sampler proposed poses to simultaneously pick many objects, followed by a learned policy for executing the pick. Simultaneously grasping multiple objects is an efficient strategy because it avoids repeated hand repositioning, which reduces total execution time, but the main limitation is that objects must be spatially close and of similar size and shape. In comparison, our method can handle scattered objects of different shapes and sizes by sequentially picking one at a time. The other works that do sequential multi-object grasping [22, 25] restrict the grasping to a maximum of two objects [22] or to primitive object shapes such as cylinders or spheres [25]. In comparison, our method can handle up to four objects of complex shapes and sizes. We also use our method to collect the largest sequential multi-object grasping dataset to date.
@@ III Problem Formulation
$\operatornamewithlimits{argmin}_{\mathbf{g}_{n}}E(\mathbf{g}_{n},\mathbf{O}_{n},\mathcal{G}_{n-1},$ $\;\mathcal{O}_{n-1},\mathcal{OS}_{n}),\forall n=1,\dots,N,$ (1)
where $\mathcal{G}_{n-1}=\{\mathbf{g}_{i}\}_{i=1}^{n-1}$, $\mathcal{O}_{n-1}=\{\mathbf{O}_{i}\}_{i=1}^{n-1}$, $\mathcal{G}_{0}=\varnothing$, and $\mathcal{O}_{0}=\varnothing$. $E$ in Eq. 1 is a differentiable function that quantifies how well grasp $\mathbf{g}_{n}$ can pick object $\mathbf{O}_{n}$ with the DoF $\mathcal{OS}_{n}$ given all previously generated grasps $\mathcal{G}_{n-1}$ and objects $\mathcal{O}_{n-1}$. We assume all previously grasped objects $\mathcal{O}_{n-1}$ remain fixed in the hand when generating a grasp on the next object $\mathbf{O}_{n}$.
@@ IV-A Opposition Space Selection Strategy
The primary objective in sequential multi-object grasping is to maximize the hand’s remaining DoF after each grasp. For this purpose, we propose a grasp planning strategy guided by OSes [30, 11, 31, 25]. An OS is a functional subspace within the hand’s kinematic structure formed by pairs of opposing surfaces (such as fingertips, lateral surfaces of fingers, or palm surfaces) along with the joints that control these surfaces [25]. It represents regions where opposing forces can be applied to create stable grasps. The number of OSes are hand-dependent and vary based on the kinematic structure. Fig. 3(a) shows the seven different OSes for the Allegro Hand.
Mathematically, each opposition space can be represented as a pair $\mathcal{OS}_{i}=\{\mathbf{J}_{i},\mathbf{S}_{i}\}$, where $\mathbf{J}_{i}\in\{0,1\}^{K}$ is a binary vector indicating which joints are involved in controlling the opposition space, and $\mathbf{S}_{i}\in\mathbb{R}^{3\times M_{i}}$ represents the 3D points on the hand where opposing forces can be applied. Fig. 3(b) shows an example of two different $\mathbf{S}_{i}$ for the Allegro Hand, where palm and pad oppositions have contact points located on the inner surfaces of fingers and palm and side oppositions have contact points on the fingers’ lateral surfaces.
@@ IV-B Optimization-based Grasp Generation
We initialize $\mathbf{g}_{n}$ at a randomly sampled position on the expanded convex hull of the target object $\mathbf{O}_{n}$ as exemplified in Fig. 4(a),. If $n=1$, then $\bm{\theta}_{1}$ is set to a natural open-hand and collision-free posture, as demonstrated in Fig. 3(b), all fingers are fully extended to create sufficient space for grasping objects with any OS. While for $n\geq 2$, $\bm{\theta}_{n}=\bm{\theta}_{n-1}$. A visual example of the optimization process when grasping a second object is shown in Fig. 4(b).
@@ IV-C Energy Function
The force-closure term ($E_{\text{fc}}$) encourages the grasp to be in force-closure equilibrium [33]. The definition for ($E_{\text{fc}}$) originates from [7], where zero friction and uniform contact force magnitudes are assumed because it significantly speeds up the computation. We follow the same definition and write it as
@@ V Dataset Generation
max width=0.9 OSes Success Total Success Rate (%) Middle-Ring 100.83 704.97 14.30 Index-Middle 97.43 700.90 13.90 Thumb-Index 97.29 1691.41 5.75 Ring-Palm 147.33 323.31 45.57 Middle-Palm 152.77 404.84 37.74 Index-Palm 194.13 546.60 35.52 Thumb-Palm 85.86 485.80 17.67 Num. Obj. Grasped Consumed Total Consumed Rate (%) One 0.00 444.72 0.00 Two 9.75 251.52 3.88 Three 69.20 139.31 49.68 Four 40.09 40.09 100.00
@@ VII Experiments
Refer to captionSeqDatasetGrasp’Em
Figure 5: Qualitative results. For SeqGrasp (first row) and SeqDiffuser (second row), the grasp sequences are visually indicated by a color gradient, transitioning from lighter to darker shades. In contrast, for MultiGrasp (third row), the color gradient is only used to differentiate the objects. For SeqGrasp and SeqDiffuser, we only show consumed grasps, that is, when $\mathcal{OS}=\emptyset$.
We experimentally evaluate SeqGrasp and SeqDiffuser in both simulation and the real world. The specific questions we aim to address with the experiments are:
@@
We compare our method to the optimization-based sampler MultiGrasp from [13], which generates simultaneous grasps on clustered objects. As such, for MultiGrasp, we must first sample clustered object configurations and then generate multi-object grasps directly on the object cluster. In contrast, SeqGrasp and SeqDiffuser do not require objects to be spatially close as they generate grasps sequentially based on previously successful ones. While this comparison is not entirely fair, we believe that comparing these two strategies still offers valuable insights.
@@ VII-A Simulation Experiments
    Diversity (Div.) in radians: Grasp diversity is determined by calculating the standard deviation of $\mathbf{g}$ across all successful grasps.
@@
    Efficiency (Eff.) in seconds: The computational time required to generate a batch of 256 grasps on an NVIDIA A100.
@@
max width=0.9 SR $\uparrow$ Pene. $\downarrow$ Eff. $\downarrow$ Div. $\uparrow$ Method SData G’Em SData G’Em Avg. Avg. MulG-1 66.84 65.39 1.14 1.27 600 0.284 SeqD-1 46.95 45.23 5.55 5.59 0.8 0.321 SeqG-1 50.04 40.78 1.73 2.16 900 0.332 MulG-2 22.46 16.48 2.30 2.83 750 0.347 SeqD-2 18.83 23.00 5.84 6.28 0.8 0.359 SeqG-2 21.21 32.03 2.14 1.78 900 0.367 MulG-3 10.78 3.55 3.39 4.04 900 0.340 SeqD-3 11.05 9.22 5.93 6.47 0.8 0.334 SeqG-3 19.49 21.05 2.23 2.23 900 0.349 MulG-4 0.90 0.47 5.17 6.27 1000 0.329 SeqD-4 3.01 1.68 6.18 6.69 0.8 0.293 SeqG-4 2.93 5.04 2.70 2.62 900 0.312
TABLE III: Simulation results. MulG, SeqG, SeqD, G’Em, and SData are short for the MultiGrasp, SeqGrasp, SeqDiffuser, Grasp’Em, and SeqDataset, respectively. The $-i$ following the method name denotes the number of objects used for grasp generation. $\uparrow$($\downarrow$), the higher (lower), the better.
The quantitative results are presented in Table III, while Fig. 5 qualitatively illustrates a few grasps. The results demonstrate that SeqGrasp achieves the highest success rate and the lowest penetration depth when grasping two or more objects.
Figure 6: Number of hand-object contacts for different methods.
However, the lower performance of SeqGrasp on single-object grasps is attributed to the fewer contact points compared to MultiGrasp, as shown in Fig. 6. For single-object grasps, SeqGrasp and SeqDiffuser establish approximately two contacts on average, whereas MultiGrasp achieves three or more due to it engaging all the available DoF. However, as the number of objects increases, the overall hand-object penetration increases for MultiGrasp but not for SeqGrasp.
We observe a notable decline in performance across all methods when transitioning from three-object to four-object grasps. We hypothesize that this decline occurs because the three previously grasped objects occupy substantial space within the Allegro Hand, pushing the fourth object grasp to the limits of the hand’s kinematic redundancy. Additionally, as the number of grasped objects increases, object-object interactions grow exponentially, making the task considerably more challenging, a finding also reported in [13]. Nevertheless, SeqGrasp demonstrates superior performance in scenarios involving three or more objects.
Another notable observation is that SeqDiffuser generates grasps with high penetration, which aligns with previous work on single-object diffusion-based grasp sampling [3]. Still, SeqDiffuser is valuable because it generates grasps 750-1250 times faster than SeqGrasp and MultiGrasp.
@@ VII-B Real-World Experiments
To evaluate grasp stability in the real world, we tested the generated grasps on a real Allegro Hand mounted to a Franka Emika Panda robotic arm. We evaluated the grasps on the two object sets visualized in Fig. 7: one comprising eight 3D printed SeqDataset test objects, and another comprising eight everyday objects. To enhance surface friction, we coated the printed object with silicone gel and attached an anti-slip grip to the palm.
We followed the procedure outlined in [25] to replicate the grasp on the real hardware. This procedure involves positioning each object as closely as possible with its generated pose and then closing the relevant joints of the hand to retain the objects. For grasps with severe penetration issues, we placed the object in the closest feasible real-world position. For MultiGrasp and SeqGrasp, we replicated the successful grasps with the lowest energy, while for SeqDiffuser, we randomly chose one of the generated grasps. Due to the larger real-world objects, four-object grasping exceeded the workspace limits of the Allegro Hand. Thus, we only evaluated three-object grasps on that object set. In contrast, both three- and four-object grasping were tested on the 3D-printed object set.
We evaluated the grasps with the grasp stability test shown in Fig. 7. In this test, the arm followed a randomly generated dynamic trajectory executed with the highest achievable acceleration. The arm was then moved to a holding position with the palm facing downward. From this position, the last joint rotated by $\pm$90°. This test comprehensively assessed the grasp stability by evaluating both translational and rotational disturbances.
Figure 7: Left: Two object sets for the grasp stability test. Right: The dynamic stability test procedure.
@@
max width= Real exp. SeqG-3 SeqG-4 SeqD-3 SeqD-4 MulG-3 MulG-4 Printed Obj. Succ. trials 5/10 5/10 4/10 0/10 1/10 1/7 Succ. objects 2.3$\pm$0.9/3 3.4$\pm$0.8/4 2.1$\pm$0.9/3 1.5$\pm$0.8/4 1.2$\pm$1.1/3 1.6$\pm$1.5/4 Real Obj. Succ. trials 5/10 / 0/10 / 0/6 / Succ. objects 2.1$\pm$1.2/3 / 1.8$\pm$0.4/3 / 1.0$\pm$0.6/3 /
TABLE IV: Real-world experimental results. Each entry reports the number of successful trials out of 10 and the average number of objects successfully grasped in mean±$\pm$±std after the test.
The experimental results are presented in Table IV. Similar to the simulation results, SeqGrasp outperformed SeqDiffuser and MultiGrasp in both three-object and four-object grasping tasks, achieving an average success rate of 50%. The low success rate of SeqDiffuser (13%) is primarily due to substantial object inter-penetration in the generated grasps, which meant that the closest physically feasible grasp replicable on the real hand differed substantially from the intended configuration. MultiGrasp performed worst, with an average success rate of 8.7%. The reason is its reliance on object-object contacts for grasp stability, which are sensitive to minor object displacements. This resulted in frequent failures, particularly in four-object grasping trials where MultiGrasp achieved no successful trials and failed to generate valid grasps in three of ten attempts (last column in Table IV). Accounting for these additional failed trials, the average success rate for MultiGrasp decreases to 6.7%.
The real-world experiments highlight that our sequential grasping strategy enables robust and effective dexterous multi-object manipulation. By generating grasps iteratively and without requiring spatial proximity between objects, SeqGrasp and SeqDiffuser mitigate the challenges posed by object interactions, making it a more reliable and practical approach for real-world applications.
VIII Limitations and Future work
Our work has two main limitations. First of all, neither SeqGrasp nor SeqDiffuser accounts for object sequences as OSes and object orders are uniformly sampled. The random sampling results in a diverse dataset that is useful for training and benchmarking, rather than the most optimal grasp sequence. Learning the optimal grasp sequences from human preferences is an interesting future work direction.
The second limitation is that the grasp generation process does not account for motion planning to the grasp pose. Consequently, it proved extremely challenging to find a collision-free path to the intended grasp pose, which is the primary reason no such experiments were conducted. However, we believe this limitation can be resolved by learning a policy conditioned on the final grasp pose using, e.g., reinforcement learning.